OpenStudy (anonymous):

The area of a rectangle is 18.3cm^2 and its width is 3.7cm shorter then its length. Calculate the dimensions of the rectangle correct to 3 significant figures.

jigglypuff314 (jigglypuff314):

Area of a rectangle = length * width given = 18.3 and width = length - 3.7 then 18.3 = (L)*(L - 3.7) then find L...

jigglypuff314 (jigglypuff314):

you would have to use the quadratic formula then, and so I am getting that L = -2.81 and 6.51 then plug that in from W = L - 3.7 W = -6.51 and 2.81 (rounded to 3 sig figs of course) :P

OpenStudy (anonymous):

thank you